could you show me how to do this problem 9-4m=19
Found 2 solutions by Alan3354, algebrahouse.com:
Answer by Alan3354(69443)

9-4m=19
--------
Subtract 9 from both sides.
Multiply by -1.
Divide by 4.

Answer by algebrahouse.com(1659)

9 - 4m = 19
-4m = 10 {subtracted 9 from both sides}
m = -10/4 {divided both sides by -4}
m = -5/2 {reduced}

For more help from me, visit: www.algebrahouse.com